A discrete random variable x has an expected value of 45 and variance of 16. Let us say we double every x's within the probability distribution. What will be the new variance?

1
Expert's answer
2022-01-24T02:49:51-0500

"Var(x)=16\\Var(2x)=2^2var(x)\\
=4\times 16=64"